# Meta results on data-driven control of nonlinear systems

### v1

@@ III-B On the necessity of Assumption 1
Assume that the matrix $\begin{bmatrix}Z_{0}\\ U_{0}\end{bmatrix}$ has full row rank. Then, Problem 1 is solvable if and only if Assumption 1 holds.
The above result states that for “rich” datasets, Assumption 1 is both necessary and sufficient for solving Problem 1. Next, we argue that in case the rank condition in Corollary 7 does not hold, Assumption 1 is still generically “necessary” for solving (the model-based) Problem 1 using only system data in (4). This is true providing that the closed-loop property characterized by $\mathcal{F}_{\rm des}$ is “binding”. This property is rather mild as formalized below:
@@
Note that that the binding property is independent of the data set. The property states that for any $F^{\star}\in\mathcal{F}_{\rm des}$ and nonzero vector $v$, there exists a vector $w$ such that the perturbed matrix $F^{\star}+wv^{T}$ leaves the desired set $\mathcal{F}_{\rm des}$. The following lemma provides two notable special cases for binding desired sets.
@@
Therefore, choosing $\left\|w\right\|$ arbitrary large makes $F^{\star}+wv^{T}$ unbounded in Frobenius norm and thus in any matrix norm. Hence, there always exists $w$ such that, $F^{\star}+wv^{T}\notin\mathcal{F}_{\rm des}$ implying that $\mathcal{F}_{\rm des}$ is binding $\blacksquare$
@@
Interestingly, the asymptotic stability property becomes binding if some guarantees on the region of attraction are also required in $\mathcal{F}_{\rm des}$. To see this, suppose that the desired closed-loop systems are given by those whose equilibrium is asymptotically stable, and there exists a compact estimate of region of attraction [^2] An estimate of region of attraction is a forward invariant set such that any solution initialized in this set asymptotically converges to the origin. $\Omega$ with $\mathcal{B}_{\delta_{1}}\subseteq\Omega\subseteq\mathcal{B}_{\delta_{2}}$ for some prescribed $\delta_{1},\delta_{2}>0$. The lower bound is to exclude arbitrary small estimates of region of attraction, whereas the upper bound enforces compactness of $\Omega$. To investigate if this property is binding, given $v\in\mathbb{R}^{s}$, consider the family of systems
@@
The value of $|v^{T}Z(x_{0})|$ is nonzero under the mild assumption that the functions in $Z(\cdot)$ are linearly independent. Initializing (13) at $x(0)=x_{0}$, we have
@@
Clearly, as $w$ becomes arbitrarily large, $x(1)$ leaves any given compact set $\Omega\subseteq\mathcal{B}_{\delta_{2}}$, and we find that $\mathcal{F}_{\rm des}$ is binding in this case.
@@
Assume that $\mathcal{F}_{\rm des}$ is binding and $\mathcal{F}_{\rm att}^{\rm data\,}\cap\mathcal{F}_{\rm des}=\emptyset$. [^4] The empty intersection means that Assumption 1 does not hold. Suppose that there exists a feedback $u=KZ(x)$ solving Problem 1, i.e., $A+BK\in\mathcal{F}_{\rm des}$. Then, there exist $\bar{A}$ and $\bar{B}$ consistent with the data, namely $X_{1}=\bar{A}Z_{0}+\bar{B}U_{0}$ such that $\bar{A}+\bar{B}K\notin\mathcal{F}_{\rm des}.$
Proof. By the hypotheses of the theorem, we find that $A+BK\notin\mathcal{F}_{\rm att}^{\rm data\,}$, i.e,
$\operatorname{im}\begin{bmatrix}A+BK\\ I_{s}\end{bmatrix}\nsubseteq\operatorname{im}\begin{bmatrix}X_{1}\\ Z_{0}\end{bmatrix}.$
Hence,
@@ III-C Extension to open systems
In this subsection, we show how the results can be extended to the control law $u=KZ(x)+K_{r}r$, with control gains $K\in\mathbb{R}^{m\times s}$, $K_{r}\in\mathbb{R}^{m\times m_{r}}$ and an external input $r\in\mathbb{R}^{m_{r}}$ with $m_{r}\leq m$. This results in the open system
@@
Theorem 11 provides a meta-procedure for solving Problem 2 using data:
@@
    Choose $K\in\mathcal{K}$.
Example 12 (Model reference control)
@@
Now, assume that the functions in $\bar{Z}$ is a subset of those in $Z$ [^5] If this is not the case, the library $Z(\cdot)$ can be simply extended to include any additional functions appearing in the reference model. and partition $Z$ as $Z(x)=\operatorname{col}(\bar{Z}(x),\tilde{Z}(x))$, with $\tilde{Z}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{s-\bar{s}}$. Then, the MRC problem can be equivalently recast as finding the matrices $K$ and $K_{r}$ such that
@@
In the special case where the functions in $Z(\cdot)$ coincides with those in the reference model, the control law reduces to $u=U_{0}G_{11}\bar{Z}(x)+U_{0}G_{2}r.$ A notable instance of such a special case is MRC in linear systems where $Z(x)=\bar{Z}(x)=x$. For linear systems, the resulting controller gains coicide with those reported in [19, 20, 21].
@@
The above dissipation inequality holds if and only if [22], [23, Prop. 4. 1. 2],
$\frac{\partial S}{\partial x}^{T}(x)(A+BK)Z(x)\leq 0$ (22)
and
$\frac{\partial S}{\partial x}^{T}BK_{r}=h^{T}(x).$ (23)
Suppose that the gradient of the storage function can be written as a linear combination of the functions in the library, i.e. $\frac{\partial S}{\partial x}=MZ(x)$ fo some matrix $M\in\mathbb{R}^{n\times s}.$ Similarly, we write the output as $h(x)=NZ(x)$ with $N\in\mathbb{R}^{m_{r}\times s}.$ The conditions (22)-(23) are then satisfied if
@@
:=$\mathcal{M}:=\{M\mid M\frac{\partial Z}{\partial x}=\frac{\partial Z}{\partial x}^{T}M^{T},\quad\forall x\}.$ (26)
@@
$\operatorname{im}\begin{bmatrix}F^{\star}&F_{r}^{\star}\\ I_{s}&0\\ K&K_{r}\end{bmatrix}\subseteq\operatorname{im}\begin{bmatrix}X_{1}\\ Z_{0}\\ U_{0}\end{bmatrix}$
[…]
Note that the above program does not contain any model information apart from the library $Z(\cdot)$. By tweaking the above program and putting in some additional effort, we obtain the following result as a special case.
@@
Then the control law $u=Kx+K_{r}r$ with $K=U_{0}G_{1}$ and $K_{r}=U_{0}G_{2}$, renders the controlled system (15) passive from input $r$ to output $y:=G_{2}^{T}X_{1}^{T}MZ(x)$. Moreover, a storage function certifying passivity is given by
[…]
@@
We note that after applying the control law stated in Corollary 14, the resulting controlled system (15) takes the form
@@
By using (29c), it turns out that the above controlled system admits a port-Hamiltonian representation with $S$ serving as the Hamiltonian [23, Ch. 6]. The fact that port-Hamiltonian systems are passive proves the claim made in Corollary 14; namely,
@@
where we used (29c) to write the inequality. The explicit form of the storage function in (30) follows from the Hadamard lemma [24, Ch. 2], noting $\frac{\partial S}{\partial x}=MZ(x)$. It is worth mentioning that a systematic way for feedback passivation of nonlinear systems is in general missing even in the model based regime. Specialising the class of passive systems to port-Hamiltonian systems allows to write the more tractable conditions (29) in place of (13). For linear systems, the two classes coincide. In particular, restricting the results to linear systems $Z(x)=x$ and positive definite storage functions, the class (26) reduces to positive definite $n\times n$ matrices and the constraints in (29) reduce to
@@ IV Conclusions
In this note, we have introduced a comprehensive, versatile, and unifying framework for direct data-driven control of nonlinear systems. This framework addresses the sparsity observed in the literature and put forward systematic procedures for synthesizing control algorithms directly from data. By adopting a meta-framework approach, we have shown how several existing results can be incorporated and extended, enhancing the application of data-driven methodologies to various control scenarios including nonlinear stabilization, nonlinear oscillator design, model reference control, and passivizing feedbacks. The examples discussed demonstrate not only the applicability of the proposed framework, but also its potential to foster further advancements in data-driven control research. As the challenges of nonlinear systems and the data they generate continue to grow, the need for effective data-driven control solutions becomes increasingly critical. While the discussed framework is well-positioned to meet these challenges, there is still a long way to go in providing complete answers to the complexities inherent in nonlinear data-driven control.

### v2

@@ III-B On the solvability of Problem 1 from data
Assume that the matrix $\left[\begin{smallmatrix}Z_{0}\\ U_{0}\end{smallmatrix}\right]$ has full row rank. Then, Problem 1 is solvable if and only if Assumption 1 holds.
The above result states that for ‘rich’ datasets, Assumption 1 is both necessary and sufficient for solving Problem 1. For linear systems, where $Z(x)=x$, this richness property, specifically the full row rank condition of $\begin{smallmatrix}Z_{0}\\ U_{0}\end{smallmatrix}]$, can be guaranteed by ensuring that the input is persistently exciting of order $n+1$, as established by Willems’ fundamental lemma [22]. For general nonlinear systems, to the best of our knowledge, a full extension of the fundamental lemma is not yet available. Nevertheless, tailored formulations have been developed for specific classes of nonlinear systems. We refer to [8, Sec. 4.1] for an overview of such results.
Next, we argue that in case the rank condition in Corollary 8 does not hold, the condition in Assumption 1 is still generally “necessary” for solving Problem 1 using only system data in (3). This is true providing that the closed-loop property characterized by $\mathcal{F}_{\rm des}$ is “binding”. This property is rather mild as formalized below:
@@
Note that that the binding property is independent of the data set. The property requires that for any $F^{\star}\in\mathcal{F}_{\rm des}$ and nonzero vector $v$, there exists a vector $w$ such that the perturbed matrix $F^{\star}+wv^{T}$ leaves the desired set $\mathcal{F}_{\rm des}$. The following lemma provides two notable special cases for binding desired sets.
@@
Therefore, choosing $\left\|w\right\|$ arbitrary large makes $F^{\star}+wv^{T}$ unbounded in Frobenius norm and thus in any matrix norm. Hence, for any $v$, there always exists $w$ such that, $F^{\star}+wv^{T}\notin\mathcal{F}_{\rm des}$ implying that $\mathcal{F}_{\rm des}$ is binding $\blacksquare$
@@
Interestingly, the asymptotic stability property becomes binding if some guarantees on the region of attraction are also required in $\mathcal{F}_{\rm des}$. To see this, suppose that the desired closed-loop systems are given by those whose equilibrium is asymptotically stable, and there exists a compact estimate of region of attraction [^2] An estimate of region of attraction is a forward invariant set such that any solution initialized in this set asymptotically converges to the origin. $\Omega$ with $\mathcal{B}_{\delta_{1}}\subseteq\Omega\subseteq\mathcal{B}_{\delta_{2}}$ for some prescribed $\delta_{1},\delta_{2}>0$. Here, $\mathcal{B}_{\delta}:=\{\left\|x\right\|\leq\delta\}$ denotes the ball of radius $\delta$ centered at the origin, the lower bound is to exclude arbitrary small estimates of region of attraction, and the upper bound enforces compactness of $\Omega$. To investigate if this property is binding, given $v\in\mathbb{R}^{s}$, consider the family of systems
@@
The value of $|v^{T}Z(x_{0})|$ is finite and nonzero under the mild assumption that the functions in $Z(\cdot)$ are linearly independent. Initializing (16) at $x(0)=x_{0}$, we have
@@
Clearly, as $w$ becomes arbitrarily large, $x(1)$ leaves any given compact set $\Omega\subseteq\mathcal{B}_{\delta_{2}}$, and thus $\mathcal{F}_{\rm des}$ is binding in this case.
@@
Assume that $\mathcal{F}_{\rm des}$ is binding and that $\mathcal{F}_{\rm att}^{\rm data\,}\cap\mathcal{F}_{\rm des}=\emptyset$. [^4] The empty intersection means that Assumption 1 does not hold. Suppose that there exists a feedback $u=KZ(x)$ solving Problem 1, i.e., $A+BK\in\mathcal{F}_{\rm des}$. Then, there exist $\bar{A}$ and $\bar{B}$ consistent with the data, namely $X_{1}=\bar{A}Z_{0}+\bar{B}U_{0}$ such that $\bar{A}+\bar{B}K\notin\mathcal{F}_{\rm des}.$
Proof. By the hypotheses of the theorem, we find that $A+BK\notin\mathcal{F}_{\rm att}^{\rm data\,}$, i.e, $\operatorname{im}\left[\begin{smallmatrix}A+BK\\ I_{s}\end{smallmatrix}\right]\nsubseteq\operatorname{im}\left[\begin{smallmatrix}X_{1}\\ Z_{0}\end{smallmatrix}\right].$ Hence,
@@ III-C Extension to open systems
In this subsection, we show how the results can be extended to feedforward-feedback control law $u=KZ(x)+K_{r}r$, with control gains $K\in\mathbb{R}^{m\times s}$, $K_{r}\in\mathbb{R}^{m\times m_{r}}$ and an external input $r\in\mathbb{R}^{m_{r}}$ with $m_{r}\leq m$. This results in the open system
@@
Theorem 12 provides a systematic procedure for solving Problem 2 using data:
@@
    Choose $\begin{bmatrix}K&K_{r}\end{bmatrix}\in\mathcal{K}_{\rm ext}$.
Example 13 (Nonlinear model reference control)
@@
Now, assume that the functions in $\bar{Z}$ is a subset of those in ZZ [^5] If this is not the case, the library $Z(\cdot)$ can be simply extended to include any additional functions appearing in the reference model. and partition $Z$ as $Z(x)=\operatorname{col}(\bar{Z}(x),\tilde{Z}(x))$, with $\tilde{Z}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{s-\bar{s}}$. Then, the MRC problem can be equivalently recast as finding the matrices $K$ and $K_{r}$ such that
@@
In the special case where the functions in $Z(\cdot)$ coincides with those in the reference model, the control law reduces to $u=U_{0}G_{11}\bar{Z}(x)+U_{0}G_{2}r.$ A notable example of this special case is MRC in linear systems where $Z(x)=\bar{Z}(x)=x$. For linear systems, the resulting controller gains coincide with those reported in [24, 25, 26].  $\square$
@@
This dissipation inequality holds if and only if [27, Prop. 4. 1. 2],
$\frac{\partial S}{\partial x}^{T}(x)(A+BK)Z(x)\leq 0$ (25a)
$\frac{\partial S}{\partial x}^{T}BK_{r}=h^{T}(x).$ (25b)
Suppose that the gradient of the storage function can be written as a linear combination of the functions in the library, i.e. $\frac{\partial S}{\partial x}=MZ(x)$ fo some matrix $M\in\mathbb{R}^{n\times s}.$ Similarly, we write the output as $h(x)=NZ(x)$ with $N\in\mathbb{R}^{m_{r}\times s}.$ The conditions (25) are then satisfied if
@@
$\mathcal{M}:=\{M\mid M\frac{\partial Z}{\partial x}=\frac{\partial Z}{\partial x}^{T}M^{T},\quad\forall x\}.$ (28)
@@
$\operatorname{im}\begin{bmatrix}F^{\star}&F_{r}^{\star}\\ I_{s}&0\\ K&K_{r}\end{bmatrix}\subseteq\operatorname{im}\begin{bmatrix}X_{1}\\ Z_{0}\\ U_{0}\end{bmatrix},\;\begin{bmatrix}F^{\star}&F^{\star}_{r}\end{bmatrix}\in\mathcal{F}_{\rm des}$ (30)
with $\mathcal{F}_{\rm des}$ given by (14). Note that the above program does not contain any model information apart from the library $Z(\cdot)$. By tweaking the above program and putting in some additional effort, we obtain the following result as a special case.
@@
Then the control law $u=KZ(x)+K_{r}r$ with $K=U_{0}G_{1}$ and $K_{r}=U_{0}G_{2}$, renders the controlled system (18) passive from input $r$ to output $y:=G_{2}^{T}X_{1}^{T}MZ(x)$. Moreover, a storage function certifying passivity is given by
$S(x)=\int_{0}^{1}x^{T}MZ(tx)dt.$ (32)
@@
We note that after applying the control law in Corollary 15, the resulting controlled system (18) takes the form
@@
Using (31c), it turns out the above controlled system admits a port-Hamiltonian representation with $S$ serving as the Hamiltonian [27, Ch. 6]. The fact that port-Hamiltonian systems are passive proves the claim made in Corollary 15; namely,
@@
where we used (31c) to write the inequality. The explicit form of the storage function in (32) follows from the Hadamard lemma [28, Ch. 2], noting $\frac{\partial S}{\partial x}=MZ(x)$. It is worth mentioning that a systematic way for feedback passivation of nonlinear systems is in general missing even in the model based regime. Specialising the class of passive systems to port-Hamiltonian systems allows to write the more tractable conditions (31) in place of (14). For linear systems, the two classes coincide. In particular, restricting the results to linear systems $Z(x)=x$ and positive definite storage functions, the class (28) modifies to positive definite $n\times n$ matrices and the constraints in (31) reduce to
@@ IV Conclusions
We have introduced a unifying and versatile framework for direct data-driven control of nonlinear systems. This framework addresses the sparsity observed in the literature and put forward systematic procedures for synthesizing control algorithms directly from data. By developing a systematic framework, we have shown how several existing results can be incorporated and extended, enabling the development of new results in the domain, particularly in scenarios like nonlinear stabilization, nonlinear oscillator design, model reference control, and passivating feedback. Remarkably, regardless of the specific instance of the control problem, the underlying data-driven design principle reduces to three steps: i) specify the set of desired closed-loop systems ($\mathcal{F}_{\rm des}$), ii) intersect this set with the set of closed-loop systems attainable by data ($\mathcal{F}_{\rm att}^{\rm data\,}$), and iii) select an element from this intersection and choose the corresponding controller gain. The examples discussed demonstrate not only the applicability of the proposed framework, but also its potential to foster further advancements in data-driven control research. As the challenges of nonlinear systems and the data they generate continue to grow, the need for effective data-driven control solutions becomes increasingly critical. While the discussed framework is well-positioned to meet these challenges, there is still a long way to go in providing complete answers to the complexities inherent in nonlinear data-driven control. A particularly relevant extension of this work would be to address the neglected nonlinearities in the considered library of functions